OpenStudy (anonymous):

Find the slope, if it exists, containing the pair of points: (-2,-8) (-15,-16)?

OpenStudy (blurbendy):

-16 + 8 = - 8 -15 + 2 = -13 slope = 8/13 Use the formula Y2 - Y1 / X2 - X1 (Remember, two negatives make a positive).